### v1

# ChartEdit: How Far Are MLLMs From Automating Chart Analysis? Evaluating MLLMs’ Capability via Chart Editing

Abstract
Although multimodal large language models (MLLMs) show promise in generating chart rendering code, chart editing presents a greater challenge. This difficulty stems from its nature as a labor-intensive task for humans that also demands MLLMs to integrate chart understanding, complex reasoning, and precise intent interpretation. While many MLLMs claim such editing capabilities, current assessments typically rely on limited case studies rather than robust evaluation methodologies, highlighting the urgent need for a comprehensive evaluation framework. In this work, we propose ChartEdit, a new high-quality benchmark designed for chart editing tasks. This benchmark comprises diverse editing instructions applied to real-world charts, with each instruction-chart instance having been manually annotated and validated for accuracy. Utilizing ChartEdit, we evaluate the performance of 10 mainstream MLLMs across two types of experiments, assessing them at both the code and chart levels. The results suggest that large-scale models can generate code to produce images that partially match the reference images. However, their ability to generate accurate edits according to the instructions remains limited. The state-of-the-art (SOTA) model achieves a score of only , highlighting significant challenges in precise modification. In contrast, small-scale models, including chart-domain models, struggle both with following editing instructions and generating overall chart images, underscoring the need for further development in this area. Code is available at https://github.com/xxlllz/ChartEdit.
ChartEdit: How Far Are MLLMs From Automating Chart Analysis? Evaluating MLLMs’ Capability via Chart Editing
Xuanle Zhao1,111Equal contribution., Xuexin Liu1,111Equal contribution., Haoyue Yang1,111Equal contribution., Xianzhen Luo2, Fanhu Zeng1, Jianling Li 3, Qi Shi4222Corresponding author., Chi Chen4222Corresponding author. 1 Institute of Automation, Chinese Academy of Sciences, Beijing, China 2 Harbin Institute of Technology, Harbin, China 3 Tianjin University, Tianjin, China 4 Tsinghua University, Beijing, China {zhaoxuanle2022, liuxuexin2022, yanghaoyue2024}@ia.ac.cn
1 Introduction
Data visualization is a crucial component in various fields, enabling individuals and organizations to present and interpret data effectively Chen et al. (2008). However, creating high-quality, visually appealing charts from scratch can be time-consuming and often requires extensive adjustments and references to documentation. This highlights the need for more efficient solutions to streamline the visualization process. Previous works have explored utilizing textual descriptions to generate visualization code automatically through the model Yang et al. (2024); Zadeh et al. (2024), which significantly reduces the time and effort required for data visualization, making it more accessible and efficient for researchers.
| Name | Output | w/ Real | w/ Cor. | Diverse | Open | Editing Instruction | ||||
| Format | Chart | Code | Types | Domain | Data | Format | Layout | Style | Text | |
| ChartCraft Yan et al. (2024) | Json | ✗ | ✗ | ✗ | ✗ | ✓ | ✓ | ✓ | ✓ | ✗ |
| Plot2Code Wu et al. (2024) | Code | ✓ | ✗ | ✓ | ✓ | ✗ | ✗ | ✗ | ✗ | ✗ |
| ChartX Xia et al. (2024) | Code | ✗ | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ✗ | ✗ |
| AcademiaChart Zhang et al. (2024c) | Code | ✓ | ✗ | ✗ | ✓ | ✗ | ✗ | ✗ | ✗ | ✗ |
| ChartMimic Shi et al. (2024) | Code | ✓ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ✗ |
| ChartEdit (ours) | Code | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
Recently, multimodal large language models (MLLMs) Liu et al. (2023a); Zhang et al. (2024a); Guo et al. (2025), which leverage the rich knowledge in large language models (LLMs) Dubey et al. (2024), have demonstrated impressive performance in processing and reasoning over various modalities, such as images Wang et al. (2024b); Zeng et al. (2024) and videos Li et al. (2024). However, much of the existing research in code generation has primarily focused on using text as the sole input Li et al. (2022); Luo et al. (2023), leaving the potential of multimodal information largely unexplored.
Although there have been recent efforts to explore multimodal code generation in specific areas like charts Shi et al. (2024); Zhang et al. (2024c) and webpages Yun et al. (2024); Si et al. (2024), these works typically address direct chart/web-to-code tasks, which are akin to using code for captioning. These approaches often lack diverse instructions and fail to account for interactive or iterative modifications. This raises an important question: Can current MLLMs function like expert chart analysts, generating modified chart code based on both the source chart and detailed editing instructions?
This is a far more complex challenge, as it requires MLLMs to not only extract relevant information from the chart but also generate corresponding code and adapt it according to the provided editing instructions. However, there is currently no high-quality, diverse benchmark for evaluating chart editing tasks. Previous works, such as ChartMimic Shi et al. (2024), focus mainly on data-centric modifications, while the ChartLlama dataset lacks real-world charts, limiting the scope and relevance of their evaluations.
To address this gap, we introduce ChartEdit, a comprehensive evaluation benchmark consisting of real-world charts sourced from Arxiv, each accompanied by manually annotated code and chart instructions with reference edited code. The instructions and code are either human-written or initially generated by LLMs, followed by manual correction and alignment to ensure accuracy. To ensure diversity, we pre-define chart types and six types of editing instructions. We also introduce two task formats, input chart w/ code or w/o code, simulating real-world scenarios that require code editing from only the charts or charts with code. Furthermore, we establish and evaluate metrics based on execution rate, code-level accuracy, and chart-level consistency to comprehensively assess MLLMs’ capabilities. These metrics measure the models’ ability to generate executable code, maintain editing precision, and ensure consistency across the generated visualizations. We conduct experiments on three types of MLLMs: proprietary, general-domain open-source and chart-domain models. The results reveal that, despite notable advancements, state-of-the-art open-source models still show performance gaps compared to GPT-4o. Additionally, current small-scale MLLMs, including chart-domain models, continue to face significant challenges in generating editing code that accurately follows the instructions. Furthermore, we also investigate the impact of Chain-of-Thought (CoT) prompting Wei et al. (2022) and analyze the models’ performance across various editing instructions and chart types.
2 Related Works
2.1 Chart-Domain MLLMs
Recently, MLLMs have achieved superior performance on many visual tasks by leveraging connectors to bridge the gap between large language models and vision encoders Liu et al. (2023a, b). As a significant image type, chart-related tasks have received much attention. Previous works utilize a two-stage method that first extracts information from the chart and then utilizes language models to process the information Liu et al. (2022). Currently, end-to-end MLLMs are utilized to solve chart-related tasks with a unified model. ChartLlama Han et al. (2023) direct finetuning based on the existing LLaVA Liu et al. (2023b). mPLUG-Owl Ye et al. (2023) and mPLUG-Owl2 Ye et al. (2024) achieve superior performance on high-resolution chart images. ChartVLM Xia et al. (2024) utilizes a discriminator to determine whether intervention from the LLMs is required for a specific query. TinyChart Zhang et al. (2024b) proposes the token merging and PoT-based reasoning strategy to improve inference efficiency and understanding capacity. ChartMoE Xu et al. (2024) utilize the Mixture-of-Expert (MoE) method to align many data formats with charts.
2.2 MLLMs For Code
Previous advancements in LLMs have significantly contributed to code-related tasks. Notable models, such as DeepSeek Coder Guo et al. (2024) and StarCoder Lozhkov et al. (2024), have demonstrated substantial progress in tasks like code generation and error fixing. However, these models typically operate in a single-modal setting, relying on textual input, which limits their ability to address the complexity and range of problems.
Multimodal code generation has recently gained significant attention. Several works, like Design2Code Si et al. (2024) and Web2Code Yun et al. (2024), focus on assessing the capacities of MLLMs in generating HTML code for web pages. RoboCodeX Mu et al. (2024) proposes a multimodal code generation framework for robot behavior synthesis. The field of chart-to-code generation has also attracted considerable attention Zhao et al. (2025), as it focuses on generating code that accurately reproduces a given chart image. This task is challenging due to the complex visual elements in charts, demanding advanced models to accurately translate them into functional code. Recent works Wu et al. (2024); Xia et al. (2024); Zhang et al. (2024c) evaluate the capacities of MLLMs in this context. Although some works like Xia et al. (2024); Shi et al. (2024) have considered this problem, the editing instructions in their works only focus on one type of modification, which is not diverse enough.
3 ChartEdit
In this section, we first introduce the chart editing task and illustrate the data collection and generation pipeline.
3.1 Task Definition
In this work, we aim to leverage MLLMs to edit charts and generate the corresponding code as instructed. The task can be summarized as follows: given a chart image from Arxiv papers, along with an editing instruction , the model is expected to generate the corresponding visualization code after applying the edits, regardless of whether the original code is provided. Thus, the task can be represented as:
where denotes the processing model. indicates that the model can either receive (Chart w/o Code) or (Chart w/ Code) as input. is the output code in Python, which utilize libraries like Matplotlib and Seaborn to plot the edited chart.
3.2 Data Construction
3.2.1 Chart Collection and Filtering
In this study, to collect real-world chart images, we begin by crawling papers published on Arxiv using web scraping tools like BeautifulSoup. To ensure the quality of the collected images, we first retrieve the IDs and comments of papers published on Arxiv, then filter the papers by selecting those whose comments include keywords such as “submit”, “accept”, “under review” and “camera ready”. After this filtering step, we use the ArXiv API to download the LaTeX source files for the selected papers. After downloading the source files, we remove irrelevant files based on their suffixes, retaining only those ending in .png, .pdf, .jpg, and .svg. However, many of these images are not charts that can be reproduced by visualization code. To filter out high-quality chart images, we employ a two-step process. First, we use an MLLM to assess whether each image is a chart through zero-shot prompting. Despite this assessment, we found that many low-quality chart images remained. To address this, we developed a scoring mechanism, instructing the MLLM to evaluate the images based on four criteria: Aesthetics, Readability, Reproducibility, and Data Presentation Simplicity. Each image is assigned a total score of , and we filter out those scoring below . After completing this two-step filtering process, we are left with more than high-quality chart images.
3.2.2 Code Annotation
While we have collected a sufficient number of chart images, the charts crawled from Arxiv do not include the code required for reproduction. To refine our dataset and facilitate the extraction of edited code, we manually filter and write Python code that can reproduce the corresponding charts. However, some chart types are still missing. To address this, we supplement the dataset with additional charts from other platforms, such as Kaggle and the Matplotlib gallery Hunter (2007). To prevent data leakage, we manually modify the code to generate visually distinct versions of the original charts. As a result, we have successfully obtained charts, along with their corresponding code, forming <chart, code> source pairs.
3.2.3 Instruction Generation
After obtaining the chart source code, we propose two approaches for constructing editing instructions: (1) LLM-based generation and (2) human-written instructions. In the LLM-based approach, we predefine five key editing categories: style, format, layout, data, and text, each with multiple specific subtypes. By providing the LLM with both the source code and the chart image, we prompt it to generate editing instructions and the corresponding edited code. To ensure a diverse set of outputs and minimize errors, we generate at least three variations for each chart and instruction type. To ensure the instructions and code are rigorous and aligned with human expectations, we manually verify and modify all instructions, code, and edited images for consistency. This process results in a total of <chart, instruction, code> triplets. While LLM-based generation can produce a large volume of triplets, the diversity of descriptions and their relevance to real-world needs may be constrained by inherent biases in LLM outputs. To mitigate this limitation, we manually write one editing instruction for each chart with reference code, ensuring that the dataset more closely aligns with real-world requirements.
3.3 Dataset Statistics and Analysis
Using LLM generation and manual annotation, we construct our evaluation benchmark, ChartEdit, which comprises <chart, code> pairs and <chart, instruction, code> triplets. We further analyze the diversity of ChartEdit from both chart and instruction perspectives. To ensure a diverse selection of chart types, we have predefined distinct chart categories, carefully curating instances for each type. The distribution of these chart types is provided in the Section A.2. Additionally, we enhance the diversity of editing instructions by classifying them into six categories: five generated by LLMs, as outlined in Section 3.2.3, and one written by humans. The proportions of each editing category are shown in Figure 3. In addition to exploring the diversity of editing types, we also examine the different formats in which these instructions are presented, including plain text descriptions, code symbols such as marker=’*’, and hexadecimal color codes like color=’#ABCDEF’. We embed these editing instructions and compare them with instructions from other chart-to-code benchmarks using dimensionality reduction techniques. The result in Figure 4 shows embedding after dimension reduction, which indicates that different editing types are generally separated, and the human-written instructions almost cover all the instruction types.
4 Experiments
4.1 Baselines
To facilitate a more effective comparison of existing MLLMs in chart editing tasks, we benchmark three categories of widely used MLLMs: proprietary, open-source general-domain, and chart-domain models. (1) Proprietary Models: We evaluate two top-performing models in the multimodal domain: GPT-4o OpenAI (2024) and Claude-3.5-Sonnet Anthropic (2024) which represent the cutting edge in the multimodal domain. (2) Open-Source General-Domain Models: We select five competitive open-source models from a variety of sizes, listed in descending order of total parameters. These models include InternVL-V2.5-78B Chen et al. (2024), Qwen2-VL-72B Wang et al. (2024a), LLaVA Liu et al. (2023b), InternVL-V2.5-8B and Phi-3.5-Vision Abdin et al. (2024). These models represent the base models commonly used in the multimodal domain. (3) Chart-Domain Models: For this category, we focus on models already designed for chart-to-code generation tasks, including ChartLlama Han et al. (2023), TinyChart Zhang et al. (2024b), and ChartMoE Xu et al. (2024).
For the evaluation, we use the direct instruction prompting method across all models. Detailed prompts and implementation steps are provided in the Figure 11.
| Statistic | Value |
| Charts | |
| Total Chart | 233 |
| Types | 19 |
| Average size (px) | 876575 |
| Maximum size (px) | 30001600 |
| Instruction | |
| Total Instruction | 1405 |
| Edit Types | 6 |
| Code | |
| Average Length | 650 |
| Maximum Length | 3310 |
4.2 Evaluation Metrics
Given the lack of established evaluation metrics for assessing the completeness of image editing, we draw inspiration from recent research that uses LLMs as evaluators Gu et al. (2024); Zheng et al. (2024). In this work, we leverage GPT-4o to evaluate two key aspects: first, whether the generated code aligns with the provided editing instructions, and second, whether the generated code effectively produces the intended chart.
| Model | Chart w/o Code | Chart w/ Code | ||||
| Exec.Rate | Code-Level | Chart-Level | Exec.Rate | Code-Level | Chart-Level | |
| Proprietary Models | ||||||
| GPT-4o OpenAI (2024) | 91.46 | 59.96 | 79.87 | 98.89 | 89.96 | 93.68 |
| Claude-3.5-Sonnet Anthropic (2024) | 88.22 | 47.32 | 54.92 | 89.50 | 88.99 | 81.68 |
| Open-Source General-Domain Models | ||||||
| InternVL2.5-78B Chen et al. (2024) | 79.66 | 55.67 | 70.77 | 94.31 | 92.56 | 92.81 |
| Qwen2-VL-72B Wang et al. (2024a) | 81.65 | 46.67 | 64.06 | 86.09 | 90.30 | 79.51 |
| InternVL2.5-8B Chen et al. (2024) | 62.37 | 39.24 | 45.85 | 85.20 | 90.16 | 87.36 |
| Phi3.5-Vision Abdin et al. (2024) | 67.13 | 40.19 | 37.65 | 74.44 | 82.31 | 89.16 |
| LLaVA-13B Liu et al. (2023b) | 48.75 | 11.35 | 16.71 | 30.44 | 71.88 | 25.82 |
| Chart-Domain Models | ||||||
| ChartMoE Xu et al. (2024) | 53.44 | 21.60 | 34.73 | 81.89 | 84.87 | 89.31 |
| ChartLlama Han et al. (2023) | 52.30 | 15.82 | 27.42 | 46.34 | 50.18 | 47.00 |
| TinyChart Zhang et al. (2024b) | 36.13 | 18.34 | 25.09 | 3.51 | 18.40 | 2.54 |
Therefore, we propose evaluating both the code level and the chart level. The code-level evaluation is based on the source code, editing instructions, and output code. We use GPT-4o to score the models on two aspects: Modification Accuracy and Code Completeness. Building on related approaches Shi et al. (2024); Zhang et al. (2024c), we also evaluate the model’s overall generation capabilities at the chart level. This evaluation is performed by comparing how closely the generated chart reproduces the reference edited chart (manually created). The results reflect the degree of completeness of the generated chart image. Following Shi et al. (2024), we also utilize GPT-4o to evaluate six aspects: Types, Layout, Text, Data, Style, and Clarity. Detailed information is provided in the Figure 13.
| Model | Types | Layout | Text | Data | Style | Clarity |
| GPT-4o | 18.35 | 9.49 | 15.26 | 14.87 | 13.55 | 8.53 |
InternVL2.5-78B |
16.25 | 9.03 | 13.13 | 13.59 | 11.12 | 7.91 |
InternVL2.5-8B |
10.22 | 7.68 | 8.21 | 7.21 | 6.19 | 6.34 |
| TinyChart | 4.99 | 4.60 | 4.39 | 3.71 | 3.41 | 4.16 |
| Model | Modification Acc | Code Complete |
| GPT-4o | 30.67 | 29.29 |
| InternVL2.5-78B | 29.05 | 26.62 |
| InternVL2.5-8B | 20.16 | 19.08 |
| TinyChart | 6.46 | 11.88 |
4.3 Main Results
The main results for all the MLLMs are presented in Table 3. Although open-source models have made significant strides, and in some cases even outperformed GPT-4 in various tasks Masry et al. (2022); Zhang et al. (2024a), there is still a noticeable gap when it comes to handling complex multimodal tasks Wang et al. (2024c); Shi et al. (2024). In our experiments, GPT-4o delivers the best performance on the Chart w/ Code task, achieving the highest Execution Rate, Code-Level, and Chart-Level scores. However, the Code-Level score is still not high enough, indicating that GPT-4o faces challenges with precise editing. Among open-source MLLMs, InternVL2.5-78B achieves the best performance, but it still lags behind GPT in terms of code execution rate and chart-level metrics. Our analysis shows that the largest gap between proprietary and open-source models lies in generating the complete code corresponding to an image. There is little difference in instruction-following ability between the two, and both struggle with detailed modifications in chart editing. The chart-domain models perform worse in this task, and we believe this is due to the current limitations of MLLMs in the chart-domain, particularly in handling editing tasks. Although TinyChart Zhang et al. (2024b) and ChartLlama Han et al. (2023) have been fine-tuned on relevant datasets, both lack the ability to follow editing instructions effectively. The instructions in chart editing are much more diverse than those in direct chart-to-code generation tasks.
In the Chart w/ Code task, providing the source code of the input chart leads to significant performance improvements across all proprietary and open-source general-domain models. These models can generate more accurate code, which we believe is due to the source code supplying essential information, such as the chart’s data, text, and style. This significantly alleviates the challenge for the model in extracting relevant information from the chart itself. In this context, the task becomes more like editing the code with the chart images serving as auxiliary context, which is much easier than the Chart w/o Code task. However, we observe that most Chart-Domain models perform significantly worse in this setting. Upon analyzing TinyChart, we found that it lacks training data for relevant tasks, especially those involving longer code snippets in the input. ChartLlama claims to support chart editing, but the diversity of chart types and editing instructions in its training data is limited. As a result, its performance decreases slightly when faced with inputs that include code. Overall, tasks with code are generally easier than those without, denoting chart-domain models still require substantial improvements in handling diverse instructions.
The detailed scores of chart-level are presented in Table 4, and we select four models for analysis. Since the Chart w/o Code task is significantly more challenging, we only provide detailed scores for this task. The result shows that the performance gaps between proprietary and open-domain models are most noticeable in the text, data, and style aspects, which are key pieces of information in the chart. Table 5 presents the detailed scores at the code level, highlighting that there remains a gap between the state-of-the-art open-source models and proprietary models in generating complete code. At the same time, the large-scale model has a significant advantage over the small-scale model in terms of modification accuracy.
5 Discussion
In this section, we analyze the following questions.
RQ1: Is Chain-of-Thought (CoT) prompting useful for MLLMs in chart editing tasks for proprietary models? Answer: No. In the CoT setting, we provide MLLMs with zero-shot prompting, instructing them to first understand the source chart, then analyze the editing instructions, and finally generate the code. However, the result in Figure 5 shows whether using proprietary models or open-source models, CoT prompting shows almost no improvement in performance. We analyze the thought process of GPT-4o and find that while CoT helps the model better understand the components in the chart and the editing instructions, it does not result in more accurate code generation. The detailed CoT prompts are provided in the Figure 11.
RQ2: How does the performance of MLLMs vary across different types of editing instructions? In Figure 6, we compare the performance across different types of editing tasks and find that models are much more effective at generating code to modify chart types. Our analysis reveals that unlike tasks requiring first grounding and then modification of specific elements, the format conversion task focuses on altering the overall visual representation of the chart, which makes it easier to edit the code. However, at the chart level, model performance on format conversion tasks is limited. This suggests that while models can change the chart type effectively, they still struggle with more complex modifications involving other elements and fail to capture all the details of the chart.
RQ3: How does the performance of MLLMs vary across different types of charts? Figure 7 illustrates model performance across different chart types. The results show that the most significant performance discrepancy between open-source and proprietary models occurs with venn diagrams. Additionally, noticeable differences emerge in handling more complex chart types, such as errorbar and multi charts. Upon further analysis, we found that the comparable performance of InternVL2.5-7B and GPT-4o on the code-level tasks is due to the fact that this metric primarily evaluates the models’ capacity to follow instructions. Since most editing instructions for venn diagrams do not require altering the chart type itself, the performance gap at the code level is smaller than at the chart level.
RQ4: How do the error types of generated code vary across MLLMs? Figure 9 shows the types of runtime errors generated by different models. The results reveal that ValueError constitutes the largest proportion of errors, accounting for more than half of the total. Interestingly, the proportion of errors excluding ValueError tends to decrease as the model’s performance improves. Our analysis suggests that as code generation capabilities improve, many common errors are effectively mitigated while more complex issues related to data processing persist.
6 Conclusion
In this work, we introduce ChartEdit, a high-quality benchmark that includes various types of editing instructions, each of which is either manually written or first generated by an LLM and then manually corrected. We evaluate performance across proprietary, open-source general-domain, and chart-domain models. The results demonstrate that proprietary models consistently outperform others in both following editing instructions and generating accurate chart images. Additionally, we observe that current chart-domain models generally struggle with complex instructions and handling diverse inputs. We believe that automated chart editing in academic research holds great promise, but MLLMs need to further improve their ability to effectively process chart images.
Limitation
From our perspective, our work has several limitations: (1) Our works only consider textual prompts such as Chain-of-Thought. However, more visual prompt methods could be evaluated. (2) The dataset size may not be enough in some situations. Maybe a much larger evaluation benchmark could help to find more interesting findings. (3) More accurate evaluation methods. Our code-level and chart-level metrics are evaluated via LLM, so a better calculation method could be proposed.
Ethical Statement
Our research employs publicly available models and data with proper citations. This approach minimizes the risk of generating toxic content, leveraging the widely used and non-toxic nature of our datasets and prompts.
Appendix A Appendix
A.1 Instruction Generation Details
In data generation, we use InternVL2.5-78B to evaluate whether the images crawled from Arxiv are charts and to score the quality of these chart images.
A.2 Chart Statictic
We present the count of each chart type in Table 6. The largest proportions are observed in bar and line charts, which are the most common chart types in Arxiv papers.
| Chart Type | Count | Chart Type | Count |
| line | 62 | bar | 41 |
| multi | 22 | heatmap | 21 |
| radar | 13 | scatter | 12 |
| errorbar | 10 | pie | 10 |
| violin | 7 | combination | 6 |
| box | 5 | errorpoint | 4 |
| hist | 4 | 3d | 3 |
| inset | 3 | multidiff | 3 |
| hexbin | 2 | stack | 2 |
| venn | 2 |
A.3 Prompts
A.4 Human Eval
We have evaluate four popular models on Chart w/o Code setting. This results are list below
| Model | Editing | Other | Sum |
| GPT-4o | 3.88 | 3.07 | 6.95 |
| InternVL2.5-72B | 3.08 | 2.36 | 5.44 |
| InternVL2.5-8B | 1.62 | 1.25 | 2.87 |
| ChartMoE | 1.21 | 1.03 | 2.24 |